### v2

# From Heegaard diagrams to surgery

Abstract.
The precise steps of a procedure of going from Heegaard diagrams to framed link diagrams are introduced in this note.
Mathematics Subject Classification (2020): 57-01, 57K30
Keywords : 3-manifolds, knots, links, Kirby’s calculus
1. Introduction
By a surface we mean here a closed, connected and orientable surface and a manifold is a closed oriented three-dimensional manifold. Such manifolds are presentable in terms of diagrams in several ways, and we mention here the presentation in terms of Heegaard diagrams and in terms of framed links in $S^{3}$. The latter is usually called surgery presentation. The aim of this note is to provide a precise steps in transforming planar Heegaard diagrams into surgery. The details of a reverse transformation are given, for example, in [1].
We are aware that this procedure is considered as a folklore and that it could be traced back to Lickorish’s results from [7]. However, when we were faced with a concrete problem of finding surgery presentation for some small covers (see [3] or [2] for this notion), a precise algorithm for doing this was not reachable in the literature. A companion to the algorithm provided in this note is a simple example of a small cover obtained by an identification of the facets of eight tetrahedra. This identification gives immediately a (demanding) Heegaard diagram of the obtained manifold (the real projective space), and this diagram could help one to follow the steps of the algorithm. Besides the precision in defining curves and Dehn twists, we ought to point out some heuristics that could considerably shorten the procedure of transforming Heegaard diagrams into surgery and to produce a result of lower complexity.
2. Heegaard diagrams, surgery and Dehn twists
We start by fixing the terminology and the notation.
Definition 2.1.
Let $\Gamma$ be a surface of genus $g$. A system of attaching circles for $\Gamma$ is a set $\{\gamma_{1},\ldots,\gamma_{g}\}$ of simple closed curves in this surface such that:
  1. (1)
    the curves $\gamma_{i}$ are mutually disjoint,
  2. (2)
    $\Gamma-\gamma_{1}-\ldots-\gamma_{g}$ is connected.
Let $\alpha$ and $\beta$ be two systems of attaching circles for $\Gamma$. The triple $(\Gamma,\alpha,\beta)$ is a Heegaard diagram of genus $g$ (see [9], [8] and [5]).
Remark 2.2.
Let $\Gamma$ and $\Delta$ be two surfaces of genus $g$. If $\{\gamma_{1},\ldots,\gamma_{g}\}$ and $\{\delta_{1},\ldots,\delta_{g}\}$ are systems of attaching circles for $\Gamma$ and $\Delta$, respectively, then there exists a homeomorphism from $\Gamma$ to $\Delta$, such that for every $1\leq i\leq g$ it maps $\gamma_{i}$ to $\delta_{i}$. If $\Gamma$ and $\Delta$ are oriented, one can choose this homeomorphism to be either orientation preserving, or orientation reversing.
Definition 2.3.
Let $H$ be a handlebody with $g$ handles. A meridional disk is a properly embedded disk in $H$ (its boundary belongs to the boundary of $H$) such that by removing its neighbourhood from $H$ one obtains a handlebody with $g-1$ handles. A complete system of meridional disks for $H$ consists of $g$ mutually disjoint disks such that by removing their neighbourhoods from $H$ one obtains a ball.
Remark 2.4.
The boundaries of a complete system of meridional disks for $H$ make a system of attaching circles for $\partial H$. We call this system meridians of $H$.
From Remarks 2.2 and 2.4 we conclude the following.
Corollary 2.5.
Let $\{\gamma_{1},\ldots,\gamma_{g}\}$ be a system of attaching circles for a connected, orientable surface $\Gamma$. Let $H$ be a handlebody with $g$ handles and $\{\mu_{1},\ldots,\mu_{g}\}$ its complete system of meridional disks. Then there exists a homeomorphism $h\colon\Gamma\stackrel{{\scriptstyle\approx\;\>}}{{\to}}\partial H$ such that for every $1\leq j\leq g$ it maps $\gamma_{j}$ to $\partial\mu_{j}$.
Definition 2.6.
Let $M$ be a manifold and let $\Gamma$, $H$ and $H^{\prime}$ be respectively a surface, and two handlebodies embedded in $M$ such that $\partial H=\partial H^{\prime}=\Gamma=H\cap H^{\prime}$ and $M=H\cup H^{\prime}$. The triple $(\Gamma,H,H^{\prime})$ is a Heegaard splitting of $M$.
Definition 2.7.
A Heegaard diagram $(\Gamma,\alpha,\beta)$ is compatible with a Heegaard splitting $(\Gamma,H,H^{\prime})$ of a manifold, when $\alpha$ consists of meridians of $H$, while $\beta$ consists of meridians of $H^{\prime}$.
Definition 2.8.
A Heegaard diagram presents a manifold, when there exists a Heegaard splitting of this manifold, which is compatible with this diagram.
The following remark is analogous to [9, IV, Section 10.1, Theorem 10.2] (see also [8, Remark after Definition 2.3]).
Remark 2.9.
Every Heegaard diagram presents a unique manifold.
Remark 2.10.
A sufficient condition for $(\Gamma,\alpha,\alpha^{\prime})$, where $\Gamma$ is of genus $g$, to present the sphere $S^{3}$ is that $\Gamma$ is a connected sum $T^{2}_{1}\sharp\ldots\sharp T^{2}_{g}$ of $g$ copies of tori, such that $\alpha_{i}$ and $\alpha^{\prime}_{i}$ are two simple closed curves in $T^{2}_{i}$ that intersect transversally in a single point. This condition is equivalent to: for $i\neq j$, $\alpha_{i}\cap\alpha^{\prime}_{j}=\emptyset$, and for every $i$, $\alpha_{i}$ intersects transversally $\alpha_{i}$ in a single point.
Definition 2.11.
Assume that $\Gamma$, $\alpha$ and $\alpha^{\prime}$ satisfy the conditions of Remark 2.10 and $H$, $H^{\prime}$ are two copies of a handlebody with $g$ handles such that $H$ is standardly embedded in $R^{3}\subset S^{3}$ and the complement of its interior with respect to $S^{3}$ is $H^{\prime}$. The identification of $\Gamma$ with $\partial H=\partial H^{\prime}$ so that $\alpha$ become meridians of $H$, while $\alpha^{\prime}$ become meridians of $H^{\prime}$ is called a proper embedding of $\Gamma$ in $S^{3}$ with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
Remark 2.12.
Let $(\Gamma,\alpha,\beta)$ be a Heegaard diagram of genus $g$ that presents a manifold $M$ and let the curves $\alpha^{\prime}$ be introduced in this diagram according to Remark 2.10. Consider a proper embedding of $\Gamma$ with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$. If $h\colon\Gamma\to\Gamma$ is an orientation preserving homeomorphism that for every $1\leq i\leq g$ maps $\alpha^{\prime}_{i}$ to a curve isotopic to $\beta_{i}$, then $M$ is homeomorphic to the manifold $((H\times\{0\})\cup(H^{\prime}\times\{1\}))/\sim$, where $\sim$ is such that for every $x\in\Gamma$ we have $(h(x),0)\sim(x,1)$.
We present Heegaard diagrams in the following form. By cutting $\Gamma$ along the curves from $\alpha$, one obtains a “sphere with $2g$ holes”. In this way every such curve appears as the boundary of two discs removed from the sphere. The curves from $\beta$ intersecting the curves from $\alpha$ are presented by paths connecting some points on the boundaries of the removed discs. In this way, a Heegaard diagram is presented as a part of the sphere (projected to the plane) and is called planar. Such a diagram should provide an unambiguous identification of the endpoints of paths belonging to $\beta$. A system of attaching curves $\alpha^{\prime}$ is introduced in this diagram according to Remark 2.10.
A surgery is a link in $S^{3}$ whose components are labeled by integers (framings). A surgery is interpreted as a manifold by removing a tubular neighbourhood (a solid torus) of each link component and sewing it back according to the corresponding framing $n$—a meridian of the solid torus is sewed back so that it winds $n$ times in the direction of the meridian and once in the direction of the longitude (see [9]).
Definition 2.13.
Let $l$ be a simple closed curve in an oriented surface $\Gamma$, and let $A$ be an annulus embedded in $\Gamma$ with $l$ as one boundary component. We define intuitively and up to isotopy the right Dehn twist $\theta_{l}\colon\Gamma\to\Gamma$ along $l$. It is the identity in the complement of $A$, and inside $A$ it takes a path transverse to the core of $A$ to a path that wraps once around $A$ turning in the right-hand direction (with respect to the orientation of $\Gamma$). It is evident that $\theta_{l}$ is a homeomorphism and its inverse $\theta^{-1}_{l}$ is called the left Dehn twist along $l$. Both left and right Dehn twists are illustrated in Figure 1.
Figure 1. The left and the right Dehn twists
Definition 2.14.
Let $l$ and $l^{\prime}$ be two oriented simple closed curves on $\Gamma$ that intersect transversally. The sign of an intersection point of $l$ with $l^{\prime}$ is +1 when a pair of tangent vectors of $l$ and $l^{\prime}$ in this order is consistent with an oriented basis for $\Gamma$, otherwise the sign is -1. The algebraic intersection number $i_{A}(l,l^{\prime})$ of $l$ with $l^{\prime}$ is the sum of signs of intersection points of $l$ with $l^{\prime}$.
3. The algorithm
Let $M$ be presented by a Heegaard diagram $(\Gamma,\alpha,\beta)$ in which a family of curves $\alpha^{\prime}$ is introduced according to Remark 2.10, so that $(\Gamma,\alpha,\alpha^{\prime})$ presents $S^{3}$. The first step in transforming a Heegaard diagram into surgery consists in finding a composition of Dehn twists of $\Gamma$ that maps curves $\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{g}$ to the curves isotopic to $\beta_{1},\ldots,\beta_{g}$, respectively. This could be done along the lines of the first part of the proof of [7, Theorem 1]. In order to be precise in determining right or left Dehn twists (the right twists have priority in this note) assume that $\Gamma$ is oriented so that its normal is directed towards the handlebody $H^{\prime}$.
Our algorithm for finding a composition of Dehn twists of $\Gamma$ that maps curves $\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{g}$ to the curves isotopic to $\beta_{1},\ldots,\beta_{g}$ goes as follows.
Auxiliary steps ($p$ and $q$ are simple closed curves in $\Gamma$ and for (D2b) and (D3) we assume that $p$ is oriented in an arbitrary way):
  1. (D1)
    If $p$ intersects $q$ in a single point, then $\theta_{p}\circ\theta_{q}$ maps $p$ to a curve isotopic to $q$. (See [7, Lemma 1].)
  2. (D2a)
    If $q$ belongs to a system of attaching circles and $p$ intersects $q$ twice, with zero algebraic intersection, then there is an isotopy moving $p$ away from $q$ and leaving all the other curves from the system fixed. (See the second paragraph of the proof of [7, Lemma 3].)
  3. (D2b)
    Let $A$ and $B$ be two intersection points of $p$ and $q$, which are adjacent on $q$. We denote by $[A,B]$ the segment on $q$ having no other common points with $p$. Assume that $p$ is oriented in the same direction at $A$ and $B$ with respect to $q$ and by walking along $p$ in its direction one has to turn left at $A$ in order to continue the walk along $[A,B]$. Then consider the curve $p_{1}$ that consists of the part of $p$ incoming to $A$, the segment $[A,B]$, and the part of $p$ outgoing from $B$. (See [7, Lemma 2, Case 1].) The curve $\theta_{p_{1}}[p]$ is isotopic to a curve intersecting $q$ in less points than $p$.
  4. (D3)
    Let $A$, $B$ and $C$ be three intersection points of $p$ and $q$, which are adjacent on $q$ and the direction of $p$ induces the circular order $ABC$. We denote by $[A,C]$ the segment on $q$ containing $B$. Assume that $p$ is oriented in alternating directions at $A$, $B$ and $C$ with respect to $q$ and by walking along $p$ in its direction one has to turn left at $A$ in order to continue the walk along $[A,C]$. Then consider the curve $p_{1}$ that consists of the part of $p$ incoming to $A$, the segment $[A,C]$, and the part of $p$ outgoing from $C$. (See [7, Lemma 2, Case 2].) The curve $\theta_{p_{1}}[p]$ is isotopic to a curve intersecting $q$ in less points than $p$.
The isotopies mentioned in (D1), (D2b) and (D3) do not move points outside narrow neighbourhoods of $p$ and $q$, so that the curves disjoint from $p$ and $q$ remain fixed.
The algorithm:
  1. (A1)
    Suppose that for a fixed $0\leq t<g$, we have a composition $\delta$ of Dehn twists such that for $1\leq i\leq t$, $\delta[\alpha^{\prime}_{i}]$ is isotopic to $\beta_{i}$.
  2. (A2)
    Find a curve $\beta_{t+1}^{\prime}=\delta[\alpha^{\prime}_{t+1}]$. Since $\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{t},\alpha^{\prime}_{t+1}$ are mutually disjoint, $\beta_{t+1}^{\prime}$ is disjoint from $\beta_{1},\ldots,\beta_{t}$.
  3. (A3)
    If $\beta_{t+1}^{\prime}$ is isotopic to $\beta_{t+1}$, then we go back to (A1) with $t=t+1$ and $\delta=\delta$.
  4. (A4)
    If $\beta_{t+1}^{\prime}$ does not intersect and is not isotopic to $\beta_{t+1}$, then try to find a curve $l$, or two curves $l$ and $m$, disjoint from all $\beta_{i}$, for $1\leq i\leq t$, such that in the sequence $\beta_{t+1}^{\prime},l,\beta_{t+1}$ or $\beta_{t+1}^{\prime},l,m,\beta_{t+1}$ the neighbours intersect in just one point and there are no other intersection points. Then iterating (D1), one finds a composition $\delta^{\prime}$ of Dehn twists that maps $\beta_{t+1}^{\prime}$ to a curve isotopic to $\beta_{t+1}$, leaving all $\beta_{i}$, for $1\leq i\leq t$ fixed. We go back to (A1) with $t=t+1$ and $\delta=\delta^{\prime}\circ\delta$.
    If there are no such $l$ and $m$, then by using steps (D1), (D2a), (D2b) and (D3), one finds a composition $\delta^{\prime\prime}$ of Dehn twists mapping $\beta_{t+1}^{\prime}$ into a curve isotopic to one that intersects less $\beta$-curves than $\beta_{t+1}^{\prime}$. Denote this curve by $\beta_{t+1}^{\prime\prime}$. Moreover, $\delta^{\prime\prime}$ leaves all $\beta_{i}$, for $1\leq i\leq t$, fixed. Then go back to the beginning of (A4) with $\beta_{t+1}^{\prime}=\beta_{t+1}^{\prime\prime}$ and $\delta=\delta^{\prime\prime}\circ\delta$.
    We cannot remain forever in (A4) since the procedure from the preceding paragraph must eventually reach the point when $\beta_{t+1}^{\prime}$ intersects no $\beta$-curve. This enables one to reason as in the third paragraph of the proof of [7, Theorem 1] in order to find two curves $l$ and $m$ satisfying the conditions listed in the initial paragraph of (A4).
  5. (A5)
    If $\beta_{t+1}^{\prime}$ intersects $\beta_{t+1}$, then by using steps (D2a), (D2b) and (D3), one finds a composition $\delta^{\prime}$ of Dehn twists mapping $\beta_{t+1}^{\prime}$ into a curve isotopic to $\beta_{t+1}^{\prime\prime}$ that either does not intersect $\beta_{t+1}$ or intersects it in a single point. In the first case go back to (A3), with $\beta_{t+1}^{\prime}=\beta_{t+1}^{\prime\prime}$ and $\delta=\delta^{\prime}\circ\delta$. In the second case go back to (A1) with $t=t+1$ and $\delta=\delta^{\prime\prime}\circ\delta^{\prime}\circ\delta$, where $\delta^{\prime\prime}$ is a composition of Dehn twists obtained by (D1), which maps $\beta_{t+1}^{\prime\prime}$ to a curve isotopic to $\beta_{t+1}$. It is evident that this composition and the isotopy leave all $\beta_{i}$, for $1\leq i\leq t$ fixed.
After finding a composition $\theta_{\delta_{n}}\circ\ldots\circ\theta_{\delta_{1}}$ of Dehn twists that maps $\alpha^{\prime}$-curves onto $\beta$-curves, the next step in our procedure is to build a link in $S^{3}$ out of the curves $\delta_{1},\ldots,\delta_{n}$. This link consists of parallel copies of $\delta_{1},\ldots,\delta_{n}$ in the handlebody $H$. This is done by making a parallel copy of $\Gamma$ shallow inside $H$ (or “digging the trench”), layer by layer, so that $\delta_{n}$ is the deepest. In order to calculate the framing of each component of the link we rely on the following remark and lemma.
Remark 3.1.
Let us consider the (right) Dehn twist $\theta_{l}$ with respect to a curve $l$ on $\Gamma$ whose parallel copy in $H$ makes a component $\mu$ of the link. Let $k\in\mathbb{Z}$ be the self-linking of $l\subseteq\Gamma$, i.e. the linking number of $l$ and $\mu$, assuming they are codirected. Then the framing of $\mu$ is $k+1$. (It should be noted that this $+1$ comes from the fact that $\theta_{l}$ is a right Dehn twist with respect to the orientation of $\Gamma$ whose normal is directed outside of $H$—if one switches from right to left twists, i.e. to $\theta_{l}^{-1}$, then the framing of $\mu$ is $k-1$, but there is no need for this here.)
For the following lemma we assume that for every $1\leq j\leq g$, the curves $\alpha_{j}$ and $\alpha^{\prime}_{j}$ are oriented so that the intersection number $i_{A}(\alpha_{j},\alpha^{\prime}_{j})$ is $-1$. Moreover, the embedding of $\Gamma$ is proper with respect to $(\alpha,\alpha^{\prime},H,H^{\prime})$.
Lemma 3.2.
Let $(\Gamma,\alpha,\alpha^{\prime})$ be a Heegard diagram of genus $g$ for $S^{3}$, and $l$ be an arbitrarily oriented simple closed curve in $\Gamma$. The self-linking number of $l$ is equal to
(3.1) $\sum_{j=1}^{g}i_{A}(l,\alpha_{j})\cdot i_{A}(l,\alpha^{\prime}_{j}).$
Proof.
Note that the self-linking of $l$ is equal to the linking number of $l$ and its parallel, codirected copy in $\Gamma$. Let us assume that our planar Heegaard diagram looks like as in the picture below (which is possible according to Remark 2.10 with a help of isotopy). We can also assume that $l$ meets $\alpha$ and $\alpha^{\prime}$-curves transversely.
Note that apart of neighbourhoods of the pairs of curves $\alpha_{j}$, $\alpha^{\prime}_{j}$ there is no contribution to the self-linking number of $l$. We can observe each of these neighbourhoods separately. Take a look at the leftmost picture in Figure 2, which illustrates a neighbourhood of $\alpha_{j}$ and $\alpha^{\prime}_{j}$ and a pair of intersection points—one of $l$ and $\alpha_{j}$ and the other of $l$ and $\alpha^{\prime}_{j}$ (negative and positive intersection points are marked in black and red, respectively). It is evident that this pair of intersection points contributes by $-1$ to the sum (3.1).
Refer to caption Refer to caption Refer to caption
Figure 2.
On the other hand, this pair contributes to the self-linking of $l$ again by $-1$, which is justified by the picture in the middle of Figure 2 (parallel, codirected copies are added to the diagram as well as a “bridge” connecting two copies of the intersection point of $l$ and $\alpha_{j}$, which is “above” the diagram). The rightmost picture in Figure 2 shows that the orientation of $l$ does not affect the score. By summing over all these pairs, we get the desired result. ∎
We are ready to transform a given planar Heegaard diagram into a surgery through the following steps:
  1. (S1)
    Introduce the $\alpha^{\prime}$-curves in the diagram, as well as all the curves involved in the composition $\delta=\theta_{\delta_{n}}\circ\ldots\circ\theta_{\delta_{1}}$ obtained by the algorithm.
  2. (S2)
    Erase the bottom row of $\alpha$-curves and all the curves from the diagram not involved in $\delta$. Then introduce “bridges” for every curve intersecting $\alpha$-curves.
  3. (S3)
    By applying Lemma 3.2 and Remark 3.1 introduce the framing tied to every curve involved in $\delta$.
  4. (S4)
    Make parallel copies of the curves involved in $\delta$. This is done step by step starting from $\delta_{2}$.
4. An example
The algorithm for finding an appropriate composition of Dehn twists and all the steps from above are illustrated in the following example.
Consider the manifold obtained by an identification of the facets of eight tetrahedra. Initially, the tetrahedra form an octahedron by identification of twelve pairs of facets. Then the “opposite” facets of this octahedron remain to be identified. This identification of the facets of the octahedron easily delivers the Heegaard diagram illustrated at the left-hand side of Figure 3. The $\alpha$-curves in this diagram are denoted by 1, 4, 7 and 12, while the $\beta$-curves are denoted by 3, 6, 9, 10. The $\alpha^{\prime}$-curves are introduced according to Remark 2.10, and are denoted by 2, 5, 8, 11.
Figure 3. Heegaard’s diagram and the steps (S2) and (S3)
We start the algorithm for finding $\delta$ which maps the $\alpha^{\prime}$-curves into the $\beta$-curves with $t=0$ and $\delta$ the identity (the empty composition of Dehn twists). The choice of matching $\alpha^{\prime}$ and $\beta$-curves could simplify the procedure. Our choice is $(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3},\alpha^{\prime}_{4})=(2,5,8,11)$ and $(\beta_{1},\beta_{2},\beta_{3},\beta_{4})=(3,6,9,10)$. With $\beta_{1}^{\prime}=\delta[\alpha^{\prime}_{1}]=\alpha^{\prime}_{1}=2$ we go to (A4). Luckily, for the sequence of curves 2, 1, 3, the neighbours intersect in just one point and we have that the composition $\theta_{1}\circ\theta_{3}\circ\theta_{2}\circ\theta_{1}$ maps the curve $\alpha^{\prime}_{1}=2$ into a curve isotopic to $\beta_{1}=3$. Then we go to (A1) with $t=1$ and $\delta=\theta_{1}\circ\theta_{3}\circ\theta_{2}\circ\theta_{1}$.
With $\beta_{2}^{\prime}=\delta[\alpha^{\prime}_{2}]=\alpha^{\prime}_{2}=5$ we go to (A4), and again in the sequence of curves 5, 4, 6, which are disjoint from 3, the neighbours intersect in just one point and we have that the composition $\theta_{4}\circ\theta_{6}\circ\theta_{5}\circ\theta_{4}$ maps the curve $\delta[\alpha^{\prime}_{2}]=5$ into a curve isotopic to $\beta_{2}=6$. Then we go to (A1) with $t=2$ and $\delta=\theta_{4}\circ\theta_{6}\circ\theta_{5}\circ\theta_{4}\circ\theta_{1}\circ\theta_{3}\circ\theta_{2}\circ\theta_{1}$.
With $\beta_{3}^{\prime}=\delta[\alpha^{\prime}_{3}]=\alpha^{\prime}_{3}=8$ we go to (A4), and again in the sequence of curves 8, 7, 9, which are disjoint from 3 and 6, the neighbours intersect in just one point and we have that the composition $\theta_{7}\circ\theta_{9}\circ\theta_{8}\circ\theta_{7}$ maps the curve $\delta[\alpha^{\prime}_{3}]=8$ into a curve isotopic to $\beta_{3}=9$. Then we go to (A1) with $t=3$ and $\delta=\theta_{7}\circ\theta_{9}\circ\theta_{8}\circ\theta_{7}\circ\theta_{4}\circ\theta_{6}\circ\theta_{5}\circ\theta_{4}\circ\theta_{1}\circ\theta_{3}\circ\theta_{2}\circ\theta_{1}$.
With $\beta_{4}^{\prime}=\delta[\alpha^{\prime}_{4}]=\alpha^{\prime}_{4}=11$ we go to (A5) since this curve intersects $\beta_{4}=10$ in a single point. The curves 10 and 11 are disjoint from 3, 6 and 9, and the composition $\theta_{11}\circ\theta_{10}$ maps the curve $\delta[\alpha^{\prime}_{4}]=11$ into a curve isotopic to $\beta_{4}=10$. Hence, the following composition of Dehn twists maps $\alpha^{\prime}$-curves onto $\beta$-curves
$\theta_{11}\circ\theta_{10}\circ\theta_{7}\circ\theta_{9}\circ\theta_{8}\circ\theta_{7}\circ\theta_{4}\circ\theta_{6}\circ\theta_{5}\circ\theta_{4}\circ\theta_{1}\circ\theta_{3}\circ\theta_{2}\circ\theta_{1}.$
The steps (S2) and (S3) in transforming our Heegaard diagram into a surgery are illustrated at the right-hand side of Figure 3. The step (S4) is illustrated by several pictures. At the left-hand side of Figure 4 we illustrate a parallel copy in $H$ of $\delta_{2}$, which is the curve denoted by 2 in our case. It is linked with the parallel copy of the curve denoted by 1 since it is pushed deeper into $H$. Then we make a parallel copy of $\delta_{3}$, which is the curve denoted by 3 in our case (see the right-hand side of Figure 4).
Figure 4. Step (S4)
Next comes another parallel copy of the curve denoted by 1 inside a handle of $H$ (see the left-hand side of Figure 5). We proceed in the same manner in order to make parallel copies of $\delta_{5}$, $\delta_{6}$ and $\delta_{7}$, which are the curves denoted by 4, 5 and 6 in our case (see the right-hand side of Figure 5).
Figure 5. Step (S4) continued
Eventually, we obtain a surgery diagram illustrated in Figure 6. It is straightforward, by using Kirby’s calculus (see [6] and [4]), to transform this surgery into an unknot with framing 2, justifying that the manifold in question is $RP^{3}$.
Figure 6. The surgery diagram
Acknowledgements
We are grateful to Vladimir Grujić for pointing out to us the problem of surgery presentation of small covers and to Danica Kosanović for very useful comments concerning the previous version of this note. The first and the third author were supported by the Science Fund of the Republic of Serbia, Grant No. 7749891, Graphical Languages - GWORDS.